# Mitigating Heterogeneity in Federated Multimodal Learning with Biomedical Vision-Language Pre-training

### v1

@@ C.3 Backbones and Baselines
We focus on upgrading biomedical vision-language pre-training (VLP) strategies to heterogeneous federated learning scenes. Since federated heterogeneous VLP is a new topic, we adopt existing biomedical VLP baselines to federated strategies (Lu et al., 2023a). Specifically, we consider adopting simple language-image contrastive alignment (ConVIRT) (Zhang et al., 2022b; Radford et al., 2021), global-local language-image contrastive alignment(GLoRIA) (Huang et al., 2021), and Multi-Granularity Cross-modal Alignment (MGCA) (Wang et al., 2022). For federated learning strategies, we consider simple averaging (FedAvg) (McMahan et al., 2017; Lu et al., 2023a), centralized training (Local), and decentralized training (Global). For fair comparisons, we combine each federated learning strategy with different biomedical VLP methods. For baselines pre-trained in Local strategy, we report the averaged performance of the local models. For fair comparisons, we implement the method with the same backbones as our method. We utilize a ViT-Base model as our vision encoder and utilize the Bert-Base as our text encoder. We pre-process the input image following (Wang et al., 2022) to resize the radiology graph to the size of $224\times 224$ to match the vision encoder.
@@
Disease C1 C2 C3 C4 C5
@@
P.E. 17048 20865 17706 9974 10424
Consolid. 4879 2201 2221 4024 4615
Pneumoth. 14177 6962 6315 13991 7115
@@
Appendix D Additional Analysis Experiment
We report the communication costs and training costs of our method and other FL methods on both the server and client sides in Table 9. As shown in the table, the additional computational overhead introduced by our method is within an acceptable range.
Table 9: Summarization of computation costs.
Method Train Params Train time Server Cost Info. Trans.
De-centralized 172.6 M 14.62s/it 863M FLOP 172.6 M
FedAvg 172.6 M 14.62s/it 863M FLOP 172.6 M
FedMAE 226.1M 14.95s/it 1130.5M FLOP 226.1M
FedX 172.6 M 22.66s/it 863M FLOP 172.6 M
FedEMA 172.6 M 14.59s/it 863M FLOP 172.6 M
Ours [1st stage] 190.4M 15.4s/it 951.2M FLOP 190.4M
Ours [2nd stage] 190.4M 23.68s/it 89M FLOP 17.8M
Centralized 172.6 M 14.62s/it 863M FLOP 172.6 M
@@ Appendix D Additional Analysis Experiment
Our method has succesfully avoided the distortion caused by heterogeneous datasets during the local training. Empirically, we demonstrate the effectiveness of this guidance-based local training strategy. Specifically, we retrain the model on each local dataset with the same backbone methods and setting and conduct the study on the change of similarity between embeddings of paired image and text. We show the performance gap between the server and local retrained models. The reduced similarity of the image embedding and text embedding from paired data can be regarded as an indicator of ”distortion” of cross-modal alignment. As reported in Table 6 and Table 11, we observe that the average decrease on the similarity of image and text embedding of our method is much smaller compared to FedAvg baseline. Role of the number of clients that participated in the federated pre-training. We have conducted experiments on how the number of clients participating in the federated pre-training affects the performance of the learned model. As shown in the Table 16, we have observed that bringing more diverse clients for the federated pre-training is likely to improve the performance of the pre-trained model.
Appendix E Detailed Experiment Results
Here we show the detailed experiment results.
@@ Appendix E Detailed Experiment Results
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ B.3 Backbones and Baselines
We focus on enabling medical vision-language pre-training (VLP) methods to be applied to heterogeneous federated learning scenes and have considered the generalization ability of FedAID on different backbone VLP methods. Since federated heterogeneous VLP is a new topic, we adopt existing medical VLP baselines to federated strategies Lu et al. (2023a). Specifically, we consider adopting contrastive learning based methods: simple language-image contrastive alignment (ConVIRT) Zhang et al. (2022b); Radford et al. (2021), global-local language-image contrastive alignment(GLoRIA) Huang et al. (2021a), and Multi-Granularity Cross-modal Alignment (MGCA) Wang et al. (2022). Since the federated multi-modal per-training is a novel setting, we adapted FedMAE Yan et al. (2023), FedEMA Zhuang et al. (2022), FedMOON Li et al. (2021b) for comparison. These are self-supervised learning methods in the federated learning domain which also focus on tackling the data heterogeneity. For basic federated learning baselines, we consider simple averaging (FedAvg) McMahan et al. (2017); Lu et al. (2023a), decentralized training (Local), and centralized training (Global). For fair comparisons, we combine each federated learning strategy with different medical VLP methods. For baselines pre-trained in Local strategy, we report the averaged performance of the local models.
For fair comparisons, we implement the method with the same backbones as our method. We utilize a ViT-Base model as our vision encoder and utilize the Bert-Base as our text encoder. We pre-process the input image following Wang et al. (2022) to resize the radiology graph to the size of $224\times 224$ to match the vision encoder.
@@
Disease Client 1 Client 2 Client 3 Client 4 Client 5
@@
Pleural Effusion. 17048 20865 17706 9974 10424
Consolidation 4879 2201 2221 4024 4615
Pneumothorax 14177 6962 6315 13991 7115
@@
Figure 10: Role of guidance degree $\alpha$. Increasing $\alpha$ results in at most small penalty in the transferability.
Figure 11: Role of the uncertainty radius $\rho$. A small radius of $\rho$ brings higher generalization and transferability on downstream tasks.
Figure 12: Robustness & convergence on the number of com. turns.
Appendix C Additional Analysis Experiment
FedAvg Test on C1 Test on C2 Test on C3 Test on C4 Test on C5
Retrain On C1 0.092 0.078 0.08 0.073 0.082
Retrain On C2 0.054 0.049 0.046 0.047 0.048
Retrain On C3 0.069 0.072 0.07 0.066 0.067
Retrain On C4 0.092 0.09 0.084 0.093 0.088
Retrain On C5 0.073 0.08 0.07 0.068 0.087
Avg 0.076 0.0738 0.07 0.0694 0.0744
@@ Appendix C Additional Analysis Experiment
Strategy O.O.D In domain In domain Wst.
Rec.@1 Rec.@5 Rec.@1 Rec.@5 Rec.@1 Rec.@5
FedMAE 12.93 63.72 5.05 23.91 4.34 20.86
FedEMA 15.19 66.94 24.01 67.05 21.85 62.44
FedAvg 15.3 66.62 28.83 72.12 25.34 66.68
FedAID (Ours) 15.43 67.29 29.49 73.03 26.99 67.65
@@
Our method has succesfully avoided the distortion caused by heterogeneous datasets during the local training. Empirically, we demonstrate the effectiveness of this guidance-based local training strategy. Specifically, we retrain the model on each local dataset with the same backbone methods and setting and conduct the study on the change of similarity between embeddings of paired image and text. We show the performance gap between the server and local retrained models. The reduced similarity of the image embedding and text embedding from paired data can be regarded as an indicator of "distortion" of cross-modal alignment. As reported in Table 6 and Table 9, we observe that the average decrease on the similarity of image and text embedding of our method is much smaller compared to FedAvg baseline.
Role of the number of clients that participated in the federated pre-training. We have conducted experiments on how the number of clients participating in the federated pre-training affects the performance of the learned model. As shown in the Table 14, we have observed that bringing more diverse clients for the federated pre-training is likely to improve the performance of the pre-trained model.
Our method has better classification accuracies on unseen out-of-distribution multi-modal X-Ray datasets. The multi-modal retrieval accuracy in the unseen domain of our method beats the baseline by a large margin as shown in Table 8, which verifies our method’s ability to learn generalizable cross-modal alignment. Meanwhile, our FedAID has higher in-domain averaged and worst-case acc.. Therefore, improvements in the in-domain performance of FedAID do not come at the cost of decreased out-of-domain performance.
Convergence check: As shown in Figure 12 the downstream performance of the pre-trained got promoted as we increased the number of comm. turns. After 20 communications, the performance turned out to be stable which verified the convergent ability of our method. Detailed results can be seen in Table 12.
Appendix D Detailed Experiment Results
In addition to the experiment results shown in the above figures and tables, here we show the detailed points in Table 9, Table 10, Table 11, Table 15, Table 16, Table 17, Table 12, Table 13 and Table 14.
Table 10: The comparison of retrieval acc. on each client denoted as $\{C_{i}\}_{i=1}^{5}$,of centralized, FedAvg, and averaged acc. of decentralized pre-trained models using the ConVIRT backbone. We report the local models of decentralized pre-training strategy as Decentralizedi.
@@ Appendix D Detailed Experiment Results
Table 11: The performance of the server model after 25 commu. turns and the averaged performance of corresponding local models after 25 and 26 commu. turns, on each client denoted as $\{C_{i}\}_{i=1}^{5}$. We utilize the ConVIRT as the backbone. We report the models after local update on client datasets of FedAvg pre-training strategy as Locali.
@@
Table 15: The accuracy of the test set of each client. We show the performance of FedAvg pre-trained baseline and its retrained models on different client datasets. We report the models after local update on client datasets of FedAvg pre-training strategy as Locali.
@@
$\rightarrow$ shallow Local0 25 30.35 24.95 25.32 28.44 28.62 27.54$\downarrow$ 73.78 67.22 68.44 71.95 73.01 70.88$\downarrow$
$\rightarrow$ shallow Local1 25 34.25 26.4 27.27 29.66 30.20 29.44$\uparrow$ 78.30 69.75 72.33 75.07 77.42 74.57$\uparrow$
$\rightarrow$ shallow Local2 25 33.68 26.39 27.27 29.66 30.20 29.44$\uparrow$ 77.25 67.37 70.67 74.21 70.51 72.02$\uparrow$
$\rightarrow$ shallow Local3 25 27.65 18.94 19.30 25.61 24.87 24.95$\downarrow$ 72.40 64.19 64.82 70.47 71.13 68.60$\downarrow$
$\rightarrow$ shallow Local4 25 26.19 18.94 19.30 22.74 22.66 21.97$\downarrow$ 69.27 56.76 58.94 63.35 64.50 62.56$\downarrow$
Table 16: The accuracy of the model on each client. We show the acc. of centralized and FedAvg pre-trained baselines and de-centralized pre-trained models shown as Locali retrained on the union of training splits of client datasets. We fine-tune shallow layers of the de-centralized pre-trained model with the union dataset.
@@
Decentralized Local1 28.72 22.61 23.52 22.32 24.59 24.35 70.94 60.77 63.31 62.40 63.42 64.17
Decentralized Local2 17.36 19.85 18.23 17.58 17.68 18.14 52.70 56.05 55.14 54.01 53.54 54.29
Decentralized Local3 20.92 20.67 25.96 21.04 22.34 22.19 58.93 58.05 65.34 58.21 58.99 59.90
Decentralized Local4 20.86 20.07 20.69 25.57 21.12 21.66 59.11 56.88 57.82 64.70 58.69 59.44
Decentralized Local5 21.79 19.54 21.98 20.90 31.54 23.15 60.73 56.99 59.80 60.17 74.07 62.35
Decentralized Local1 17.68 14.71 15.37 18.18 14.14 16.02 56.99 49.56 51.29 54.55 55.10 53.50
Decentralized Local2 15.33 11.62 13.85 14.96 13.83 13.92 54.84 41.93 46.20 47.90 45.62 47.30
Decentralized Local3 17.41 13.08 14.12 15.07 15.39 15.01 50.40 44.21 46.37 49.72 52.53 48.65
Decentralized Local4 16.54 14.34 14.08 15.36 17.35 15.53 57.02 45.72 47.45 52.38 57.63 52.04
Decentralized Local5 21.68 14.34 14.20 15.56 19.50 17.06 57.89 48.43 50.18 52.36 61.66 54.10
@@
Table 18: Comparison between related tasks and our federated multi-modal pre-training task. We organize the related works based on our key areas of interest. (1). Heterogeneous clients. (2). Client-side pretraining. (3). Federated self-supervised learning (FSSL) to leverage unlabeled data. (4). multi-modal alignment. (5). Pre-train models that can be transferred to other downstream tasks. (6). No communication on public data or feature representations.
Task hetero. pretrain on client FSSL. align modalities transfer. no feat. commu.
Learn different client modalities Chen & Zhang (2022) $\checkmark$ $\checkmark$
Augment server with local modalityYu et al. (2023) $\checkmark$ $\checkmark$ $\checkmark$
FSSL. on unlabeled datasets Zhang et al. (2023a) $\checkmark$ $\checkmark$
FSSL with server fine-tune Yan et al. (2023) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Federated multi-modal pre-training $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Appendix E Problem Position
The federated vision-language pre-training is a more practical and general task compared to related works. Table 18 lists the most similar settings in previous works, and the difference between our task and theirs. Our federated multi-modal pre-training focuses on learning generalizable vision-language models which can be transferred to various downstream tasks, with paired unlabeled multi-modal data, in a fully decentralized way that can guarantee the data privacy required by clinical applications. We tackle the data heterogeneity problem in local training, which is a vital concern in federated learning. During client-server communication, we avoid data transmission on features of local data to ensure data privacy.
Appendix F Detailed Problem Formulation
We aim at mitigating the data heterogeneity issues in federally pre-training medical vision-language models from heterogeneous private client datasets. In this section, we will formulate the set-up of the Federated Vision-Language Pre-training task, which aims to capture better cross-modal alignment by leveraging diverse semantics from local datasets. Then we will introduce our assumption on multi-modal pre-training data and heterogeneity. Please see Appendix F for the detailed preliminary.
Federated Vision-Language Pre-training (FedVLP). We aim to federally pre-train medical vision-language models. Specifically, given $N$ clients $\{C_{i}\}_{i=1}^{N}$, FedVLP decentralizedly conducts vision-language representation learning across clients to learn local models, and aggregates them on a central server to obtain a more generalizable pre-trained model. Each client has a local yet private dataset $\{D_{i}\}_{i=1}^{N}$ that can’t be easily shared due to privacy issues. During FedVLP, different from supervised federated learning, we focus on the transferability of the aggregated model.
Employing an alternative updating approach, federated pre-training also consists of multiple communication turns between clients and the server the same as conventional federated learning (FL) scenes. For each turn $t$, client models $\{M_{i}^{t}\}_{i=1}^{N}$ completed local pre-training respectively and send model weights to server, then it receives and aggregates the weights to obtain server model $M^{t}$. $M^{t}$ will be broadcast to clients for updating local models, and complete the main part of the communication turn. In some FL methods, additional local computation and information transmissions are conducted after clients have received the aggregated model, and they are counted as a part of the communication. The federated vision-language pre-training is a more practical and general task compared to related works. Table 18 in appendix lists the most similar settings in previous works, and their differences with ours.
Formulation of Multi-Model Pre-training Data and Heterogeneity. We consider the multimodal datasets including vision and language modalities $X,Y$, consisting of respective data samples $X=\{x_{i}\}_{i=1}^{n_{1}},Y=\{y_{i}\}_{i=1}^{n_{2}},n_{1},n_{2}\in N^{+}$. We assume there exists discrete mutlti-label latent variables U that embed semantics across two modalities such that $X=H_{I}(U),Y=H_{T}(U)$, where $H_{I}(\cdot),H_{T}(\cdot)$ are assumed transformations that is not explicit to know. For example, the disease category of a given X-ray radiology-report pair is a latent variable connecting these two modalities, which determines the pathology region of the radiology image and semantics in the diagnosis report. Thus, we assume each client has paired image-text data that are implicitly correlated through the same latent variables. Therefore, local client datasets could be written as $\mathcal{D}:\{D_{i}:(H_{I}(u),H_{T}(u)),u\sim P_{i}(U)\}_{i=0}^{N}$, where u can be thought as a latent class. Here, we assume the local datasets are heterogeneous, which means the distribution of latent variables $P_{i}(U)$ is different across clients.
We formulate two distributions $\mu_{i}\in P(X_{i}),\nu_{i}\in P(Y_{i})$ following Chen et al. (2020) as $\mu_{i}=\sum_{j=1}^{n_{i}}a_{i}\delta_{T_{x}(u_{j}))_{i,j}}$ and $\nu_{i}=\sum_{j=1}^{n_{i}}b_{j}\delta_{T_{y}(u_{j})_{i,j}}$, with $\delta_{x}$ as the Dirac function centered on $x$. The weight vectors $a_{i}=\{a_{j}\}_{j=1}^{n_{i}}\in\Delta_{n_{i}}$ and $b_{i}=\{b_{j}\}_{j=1}^{n_{i}}\in\Delta_{n_{i}}$ belong to the $(n-1)$- dim simplex, where $\mu_{i}$ and $\nu_{i}$ are probability distributions. This formulation can also be applied to feature vectors or representations. For a given set of distributions $\{P_{i}\}_{i=1}^{N}$, we can form a new distribution $P^{c}=\sum_{i=1}^{N}w_{i}P_{i},\textbf{w}\in\Delta$ where $\Delta$ is a $(N-1)$-dim simplex by weighted averaging the sample vectors.
Model Architectures. Following the VLP works Radford et al. (2021); Yang et al. (2022), we formulate the VLP using a dual-model structure, where the model for each modality consists of a feature encoder module $f(\cdot)$ that outputs feature vectors given the raw input, and a feature alignment module $g(\cdot)$ that projects feature vectors to aligned representations in the shared embedding space. We denote the feature vectors on the text side as $l$, the feature vectors on the image side as $v$, and the aligned representation as $z_{l},z_{v}$ respectively. During the distributionally robust optimization process, while we utilize the distribution of latent semantics to construct an uncertainty set for optimization, we take local alignment modules and the server-aggregated encoder as conditions.
Uncertainty set for Distributionally Robust Optimization (DRO). After the first stage in local training, the server aggregates the weights of local encoders to get $f^{*}$ and sends it back to clients for learning a robust teacher alignment module $g^{*}$. We aim to optimize the alignment module on the worst-case distribution of an uncertainty set constructed by the local data distribution, which represents possible cases that the model would encounter. Specifically, a worst-case distribution is considered based on support distributions and the uncertain radius. We first consider the set of possible weighted combinations of support distributions and then enlarge the set by introducing an uncertain radius. We utilize distributions of latent semantics $P_{i}(U)$ of local clients as support distributions. Therefore, following the notation in F, the uncertainty set can be written as:
$\footnotesize Q^{c}:\{D(Q||P^{c})\leq\rho\},P^{c}=\sum_{i=1}^{N}w_{i}P_{i}(U),w\in\Delta_{N-1}$ (6)
where $N$ is the number of clients, $\Delta$ is a $(N-1)$-dim simplex, $D$ is a discrepancy measure, and $\rho$ is the uncertainty radius that measures the uncertainty degree of the unseen distributions and influences model’s generalization ability.
Appendix G Analysis of Theoretical Insights
G.1 Theoretical Analysis on Remark 1
Here we provide a theoretical analysis on Remark 1 in Sec. 4. During the local training on client c, given the model parameter $\theta$, we aim to show that optimizing $\min_{\theta}L_{\text{task}}(z^{c*}_{I},z^{c*}_{T})$ can be reformulated as optimizing $\min_{\theta}\left(L_{\text{task}}(z_{I}^{c},z_{T}^{c})+\alpha\left(\|z_{I}^{c}-z^{c*}_{I}\|_{2}^{2}+\|z_{T}^{c}-z^{c*}_{T}\|_{2}^{2}\right)\right)$.
Since $L_{\text{task}}(x,y)$ is $L$-lipschitz, we have:
$\|L_{\text{task}}(x_{1},y_{1})-L_{\text{task}}(x_{2},y_{2})\|\leq L(\|x_{1}-x_{2}\|+\|y_{1}-y_{2}\|).$ (7)
Hence, we have:
$L_{\text{task}}(z^{c*}_{I},z^{c*}_{T})$ $\leq\|L_{\text{task}}(z^{c*}_{I},z^{c*}_{T})-L_{\text{task}}(z^{c}_{I},z^{c*}_{T})\|+\|L_{\text{task}}(z^{c}_{I},z^{c*}_{T})\|$
$\leq L(\|z^{c}_{I}-z^{c*}_{T}\|+\|z^{c*}_{T}-z^{c*}_{T}\|)+\|L_{\text{task}}(z^{c}_{I},z^{c*}_{T})\|$
$=L\|z^{c}_{I}-z^{c*}_{I}\|+\|L_{\text{task}}(z^{c}_{I},z^{c*}_{T})\|$
$\leq L(\|z^{c}_{I}-z^{c*}_{I}\|+\|z^{c}_{T}-z^{c*}_{T}\|)+\|L_{\text{task}}(z^{c}_{I},z^{c}_{T})\|$
$=L(\|z^{c}_{I}-z^{c*}_{I}\|+\|z^{c}_{T}-z^{c*}_{T}\|)+L_{\text{task}}(z^{c}_{I},z^{c}_{T})$
Therefore, $L(\|z^{c}_{I}-z^{c*}_{I}\|+\|z^{c}_{T}-z^{c*}_{T}\|)+\|L_{\text{task}}(z^{c}_{I},z^{c}_{T})\|$ can be an upper bound of the original learning object $L_{\text{task}}(z^{c*}_{I},z^{c*}_{T})$. During local training, we add the L-2 regularization to guide the update of the local model. We also consider the contrastive loss to measure the distance between $(z^{c}_{T},z^{c}_{I})$ and $(z^{c*}_{T},z^{c*}_{I})$ in the ablation study. However, L-2 regularization brings better performance and matches our assumption.
G.2 Discussion on Cross-Modal Alignment
Here we illustrate why we model the cross-modal alignment as an alignment module that projects features of vision and language modalities to a shared representation space, and intends to minimize the distance between representations of the two modalities.
Following the notation in F,suppose we have features $I,T$ of vision and language modalities, and the corresponding alignment module $g_{I}(\cdot),g_{T}(\cdot)$. Inspired by Huang et al. (2021b), suppose $g_{I}(\cdot),g_{T}(\cdot)$ are $L$-Lipschitz, we formulate the alignment degree of two domains as:
$R_{\epsilon}=\mathbb{P}\left[u:\sup_{I=H_{I}(u),T=H_{T}(u)}\|g_{I}(I)-g_{T}(T)\|\geq\epsilon\right]$ (8)
For a simple case following Huang et al. (2021b), the goal of the alignment module can be written as:
$L_{\text{align}}(g_{I},g_{T})=\mathbb{E}_{x}\mathbb{E}_{I=H_{I}(u),T=H_{T}(u)}\|g_{I}(I)-g_{T}(T)\|^{2}$ (9)
Using the in Theorem 2 in Huang et al. (2021b) by replacing the encoder to our $g(\cdot)$, we have the following bound:
$R^{2}_{\epsilon}\leq\eta(\epsilon)^{2}\cdot\mathbb{E}_{u}\mathbb{E}_{I=H_{I}(u),T=H_{T}(u)\in A(x)}\|g_{I}(I)-g_{T}(T)\|^{2}=\eta(\epsilon)^{2}\cdot L_{\text{align}}(g_{I},g_{T})$ (10)
,where $\eta(\epsilon)$ is a infite function that corresponding to threshold $\epsilon$.
Therefore, we model the cross-modal alignment as the alignment module that minimizes the distance between the representations of two modalities.
Appendix H Limitation
The design of the proposed FedAID approach has taken it into serious consideration to mitigate the potential negative societal impact.. For example, to tackle the data privacy issue, FedAID has avoided the transmission of any data or latent representations during all server-client communications, while only transferring model parameters and loss, thereby minimizing the risk of privacy leakage in federated learning as much as possible. However, our model may still be subject to privacy-related issues due to the communication of model parameters, which is a common problem faced by other federated learning methods as well. There are also research studies in this line to avoid privacy leakage from model parameters.
For fairness, our method aims at optimizing the performance of the worst-case client to promote model equity across different cohorts, whichhas shown great generalizable ability on unseen data. This approach ensures that the pre-trained model not only performs well and fairly on the data from hospitals participating in the federated training but also demonstrates good performance on hospitals that did not participate in the training. Meanwhile we acknowledge that there may still exist limitations. Firstly, If we can resolve the trade-off between privacy protection and data representation communication, it may allow us to design more advanced algorithms to further enhance the performance of federated multimodal pre-training models.
Furthermore, in this work we mainly consider the multimodal medical datasets of X-Ray imageeport due to the limitation of the public datasets. However, some rare diseases and more complicated modalities in various medical applications lack paired data, which] may bring other challenges for federated multimodal pretraining.